# Non-Uniqueness of Metasurfaces for Wave Transformations

### v1

@@ III.3 A special class of reciprocal, passive metasurfaces:
As an initial example, consider a metasurface that completely absorbs incident waves when the electric field is aligned along the $x$ axis. For a surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$, the direction of the electric field is $\hat{\mathbf{e}}=\hat{\mathbf{x}}$, and the magnetic field is $\hat{\mathbf{h}}=-\hat{\mathbf{y}}$, with $|E|/|H|=\eta_{0}$ where $\eta_{0}$ is the free space impedance. In this case, $e_{1}=0$, $e_{2}=1$ and the reciprocal impedance (7) reduces to
$\mathrm{Z}_{s}=\eta_{0}\left(\hat{\mathbf{x}}\otimes\hat{\mathbf{x}}+\alpha\,\hat{\mathbf{y}}\otimes\hat{\mathbf{y}}\right).$ (14)
@@ IV Example: Polarization Conversion
We now take a slightly more complex illustration of the family of reciprocal metasurfaces described by Eq. (7), although we retain the assumption of surface uniformity: we consider a polarization converting metasurface. A normally incident plane wave polarized along the $x$ axis impinges on a surface with surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$. The desired field is reflected with some real amplitude $\rho$, with a $\theta$ change in the polarization angle away from $\hat{\mathbf{x}}$, as illustrated in Fig. 1. The surface electric field equals,
@@
In this example, the unit vectors (16) describing the field on the surface are purely real. Therefore, we can find a passive reciprocal metasurface using the results derived in Sec. III.3. The reciprocal surface impedance tensor is given by Eq. (7), with the vector components $e_{1}=\hat{\mathbf{e}}\cdot\hat{\mathbf{h}}$ and $e_{2}=\hat{\mathbf{e}}\cdot(\hat{\mathbf{n}}\times\hat{\mathbf{h}})$ equal to,
@@
Given the uniformity of the surface impedance in this case, we can analytically find the reflection from the surface defined by Eqns. (7), (16), and (IV) to verify the above predictions. The reflection amplitudes, $r_{xx}$, $r_{xy}$, $r_{yx}$, and $r_{yy}$ are defined such that the first index corresponds to the output polarization, and the second index to the input polarization. The expressions for these are given in appendix A. The power carried in each polarization channel is given by the square of these amplitudes, e.g. $R_{xx}=|r_{xx}|^{2}$ tells us the fraction of power reflected in the $x$ channel for incidence in the same channel. Fig. 3 shows both the results of a full wave simulation, and the reflection computed analytically, verifying both the reciprocity of the surface, $R_{xy}=R_{yx}$ and that incident $x$ polarized waves are reflected with the desired polarization. As shown in Fig. 3a, when $\alpha=0$, at designer incidence, the output power is conserved. Meanwhile for the secondary incident polarization, the surface reflects with increased power, even a divergent value when $\alpha=2$, shown in Fig. 4. For the surface to be passive (as in e.g. Chen et al. (2014b); Cheng et al. (2014); Feng et al. (2013); Huang et al. (2014); Yan et al. (2005); Yang et al. (2014); Zhu et al. (2013)), our constraint (18) shows that for $\rho=1$ (i.e. $100\%$ efficiency for $x$ polarized incidence), this is only possible when ${\rm Re}[\alpha]\to\infty$. Figs. 3b and 4 show this is indeed the case, with the amplification diminishing with increasing $\alpha$.
@@
Figure 4: Polarization multiplexed functionality: for a $\theta=\pi/4$ polarization rotator (the design field incident with $x$ polarization), incident $y$ polarized waves can be transformed very differently, depending on the value of the complex parameter $\alpha$ in Eq. (7) for a reciprocal surface impedance. The above shows that, while the proportional of reflected power $P_{x}$ for $x$ polarized incidence is always unity, that for orthogonal incidence, $P_{y}$ is subject to either absorption or amplification, depending on the sign of ${\rm Re}[\alpha]$, with lasing–like behaviour at $\alpha=2$. Note that the maximum output power, $P_{\rm max}$, here found through an appropriate combination of $x$ and $y$ polarizations always exhibits gain except as $|\alpha|\to\infty$. This is consistent with Fig. 3 and constraint (18), as $\rho=1$, $\alpha$ must become infinite for the surface to be passive.
@@
Figure 5: Real part of eigenvalues of $\mathrm{Z}_{s}$ when varying $\theta$ and $\alpha$: (a) Changing reflected polarization angle, $\theta$, at $\alpha=0$. As already indicated in Fig. 3, one eigenvalue always corresponds to loss and the other to gain within these metasurfaces. Note that as $\theta\rightarrow 90^{\circ}$, the eigenvalues tends towards ±$\pm\eta_{0}$ (i.e. a perfect absorber for the incident polarization and a infinite gain surface for the reflected polarization), denoted as black dashed lines. Note also that at $\theta=0$, both eigenvalues diverge. This does not indicate infinite gain, as such a large impedance mismatch with free space only leads to total reflection. (b) Eigenvalues for $100\%$ ($\rho=1$) and $50\%$ ($\rho=0.5$) efficient reflectors, as a function of $\alpha$. The dashed line shows the constraint (18) where the surface is passive, $\alpha>0.646$, above which the eigenvalues have a positive real part.
@@
When the two eigenvectors are orthogonal, this reduces to $\langle P_{n}\rangle=|c_{1}|^{2}\langle P_{1}\rangle+|c_{2}|^{2}\langle P_{2}\rangle>0$, which leads to the requirement that the eigenvalues of $\mathrm{Z}_{s}$ should have a positive real part. In the current example, the surface impedance matrix is both real and symmetric (provided the free parameter $\alpha$ is real). Therefore its eigenvectors are orthogonal and we simply require both eigenvalues $z_{s,1}$, $z_{s,2}$ to have a positive real part. Fig. 5 shows the real part of the two eigenvalues of the impedance tensor (7) both as a function of the angle of polarization rotation and the parameter $\alpha$. In Fig. 5 we see that the surface indeed becomes passive once the inequality (18) is satisfied.
V Summary and Conclusions
@@
Impenetrable metasurfaces are often characterized in terms of a scalar valued surface impedance, the value of which encodes both the phase and amplitude of the reflected field. However, as has been previously recognised, the use of a tensorial surface impedance—practically implemented via an array of anisotropic, polarizable elements—offers a much richer set of possible field transformations. Here we have explored the problem of designing tensorial metasurfaces for enacting arbitrary field transformations. We have investigated the uniqueness of the surface parameters for a given field transformation, when the metasurface is constrained to be passive, reciprocal, or both.
@@
Our first finding is that it is possible to find a reciprocal metasurface, whatever the wave transformation we wish to perform. The tensorial surface impedance is given by Eq. (7) and there are an infinite number of such surfaces, parameterized in terms of a single complex number, $\alpha$. In general this family of metasurfaces will exhibit gain and loss for some incident fields, showing that through carefully implementing loss and/or gain within the surface, we may obtain multiplexed functionality. As an example, we showed in Sec. IV that a metasurface rotating $x$ polarized fields by an angle $\theta$ could simultaneously show a variety of different functionalities for $y$–polarized radiation, e.g. infinite amplification, absorption, or rotation.
@@
Nevertheless, despite a large body of fascinating work (e.g. Fleury et al. (2014)) and recent experimental progress Tapar and Kishen (2021); Y. et al. (2022), metasurfaces with gain are much more difficult to implement than lossy ones. Our second finding is that, provided the complex parameter $\alpha$ satisfies the inequality (12), we can simultaneously ensure the metasurface is reciprocal and passive (without gain) for an arbitrary incident field. Surprisingly the surface parameters remain non–unique, and there remain an infinite family of surfaces that perform the same transformation. One important point is that the right hand side of the inequality (12) diverges when the initial design is made $100\%$ efficient (zero dissipation, all energy reflected). This actually reduces the design freedom, implying that the free parameter $\alpha$ must be taken to infinity in order that the surface be both lossless and reciprocal. Therefore, when the efficiency of the design is reduced (e.g. reducing the overall amplitude of the reflected field), we find that this increases the range of reciprocal, passive metasurfaces that can enact the transformation. This was demonstrated in the example of a polarization rotating metasurface, where Fig. 5b shows that reducing the efficiency immediately yields an increased range of passive reciprocal surfaces that would otherwise require ${\rm Re}[\alpha]\to\infty$. We have thus found that a careful choice of loss within the metasurface actually aids the design, allowing for polarization multiplexed functionality, at the price of reducing the efficiency.

### v3

@@ III.3 A special class of reciprocal, passive metasurfaces:
As an initial very simple example, consider a metasurface that completely absorbs incident waves when the electric field is aligned along the $x$ axis. For a surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$, the direction of the electric field is $\hat{\mathbf{e}}=\hat{\mathbf{x}}$, and the magnetic field is $\hat{\mathbf{h}}=-\hat{\mathbf{y}}$, with $|E|/|H|=\eta_{0}$ where $\eta_{0}$ is the free space impedance. In this case, $e_{1}=0$, $e_{2}=1$ and the reciprocal impedance (LABEL:eq:reciprocal_Z) reduces to
[…]
@@ IV Example: Polarization Conversion
As a second example, we now take a slightly more complex illustration of the family of reciprocal metasurfaces described by Eq. (LABEL:eq:reciprocal_Z), although we retain the assumption of surface uniformity: we consider a polarization converting metasurface. A normally incident plane wave polarized along the $x$ axis impinges on a surface with surface normal $\hat{\mathbf{n}}=\hat{\mathbf{z}}$. The desired field is reflected with some real amplitude $\rho$, with a $\theta$ change in the polarization angle away from $\hat{\mathbf{x}}$, as illustrated in Fig. 1. The surface electric field equals,
@@
[…]
@@
Given the uniformity of the surface impedance in this case, we can analytically find the reflection from the surface defined by Eqns. (LABEL:eq:reciprocal_Z), (16), and (IV) to verify the above predictions. The reflection amplitudes, $r_{xx}$, $r_{xy}$, $r_{yx}$, and $r_{yy}$ are defined such that the first index corresponds to the output polarization, and the second index to the input polarization. The expressions for these are given in appendix A. The power carried in each polarization channel is given by the square of these amplitudes, e.g. $R_{xx}=|r_{xx}|^{2}$ tells us the fraction of power reflected in the $x$ channel for incidence in the same channel. Fig. 3 shows both the results of a full wave simulation, and the reflection computed analytically, verifying both the reciprocity of the surface, $R_{xy}=R_{yx}$ and that incident $x$ polarized waves are reflected with the desired polarization. As shown in Fig. 3a, when $\alpha=0$, at designer incidence, the output power is conserved. Meanwhile for the secondary incident polarization, the surface reflects with increased power, even a divergent value when $\alpha=2$, shown in Fig. 4. For the surface to be passive (as in e.g. Chen et al. (2014b); Cheng et al. (2014); Feng et al. (2013); Huang et al. (2014); Yan et al. (2005); Yang et al. (2014); Zhu et al. (2013)), our constraint (18) shows that for $\rho=1$ (i.e. $100\%$ efficiency for $x$ polarized incidence), this is only possible when ${\rm Re}[\alpha]\to\infty$. Figs. 3b and 4 show this is indeed the case, with the amplification diminishing with increasing $\alpha$.
@@
Figure 4: Polarization multiplexed functionality: for a $\theta=\pi/4$ polarization rotator (the design field incident with $x$ polarization), incident $y$ polarized waves can be transformed very differently, depending on the value of the complex parameter $\alpha$ in Eq. (LABEL:eq:reciprocal_Z) for a reciprocal surface impedance. The above shows that, while the proportional of reflected power $P_{x}$ for $x$ polarized incidence is always unity, that for orthogonal incidence, $P_{y}$ is subject to either absorption or amplification, depending on the sign of ${\rm Re}[\alpha]$, with lasing–like behaviour at $\alpha=2$. Note that the maximum output power, $P_{\rm max}$, here found through an appropriate combination of $x$ and $y$ polarizations always exhibits gain except as $|\alpha|\to\infty$. This is consistent with Fig. 3 and constraint (18), as $\rho=1$, $\alpha$ must become infinite for the surface to be passive.
@@
Figure 5: Real part of $\mathrm{Z}_{s}$ eigenvalues for varying rotation angle, $\theta$ and design parameter, $\alpha$: (a) Changing reflected polarization angle, $\theta$, at $\alpha=0$. As already indicated in Fig. 3, one eigenvalue always corresponds to loss and the other to gain within these metasurfaces. Note that as $\theta\rightarrow 90^{\circ}$, the eigenvalues tends towards $\pm\eta_{0}$ (i.e. a perfect absorber for the incident polarization and a infinite gain surface for the reflected polarization), denoted as black dashed lines. Note also that at $\theta=0$, both eigenvalues diverge. This does not indicate infinite gain, as such a large impedance mismatch with free space only leads to total reflection. (b) Eigenvalues for $100\%$ ($\rho=1$) and $50\%$ ($\rho=0.5$) efficient reflectors, as a function of $\alpha$. The dashed line shows the constraint (18) where the surface is passive, $\alpha>0.646$, above which the eigenvalues have a positive real part.
@@
When the two eigenvectors are orthogonal, this reduces to $\langle P_{n}\rangle=|c_{1}|^{2}\langle P_{1}\rangle+|c_{2}|^{2}\langle P_{2}\rangle>0$, which leads to the requirement that the eigenvalues of $\mathrm{Z}_{s}$ should have a positive real part. In the current example, the surface impedance matrix is both real and symmetric (provided the free parameter $\alpha$ is real). Therefore its eigenvectors are orthogonal and we simply require both eigenvalues $z_{s,1}$, $z_{s,2}$ to have a positive real part. Fig. 5 shows the real part of the two eigenvalues of the impedance tensor (LABEL:eq:reciprocal_Z) both as a function of the angle of polarization rotation and the parameter $\alpha$. In Fig. 5 we see that the surface indeed becomes passive once the inequality (18) is satisfied.
V Example: Conversion of a cylindrical TE–polarized wave into a TM–polarized plane wave
In the previous section we showed the simplest possible example: a normal incidence plane wave source. We now demonstrate the applicability of our method to more general, inhomogeneous wave transformations. We use it to calculate the impedance profile required to convert an incident cylindrical wave into a plane wave. To calculate the full tensorial impedance required at the surface, we numerically modelled a line curent in free space (Transverse Electric (TE) polarization) at a distance of $\lambda/4$ from the interface and tabulated the surface field as shown in appendix B. The reflected field was chosen to be a Transverse Magnetic (TM) polarised plane wave. The sum of these two surface fields were then used as the fields in our design formula, Eq.(LABEL:eq:reciprocal_Z). The components of the resulting impedance tensor, $\bm{Z}$ are shown in Fig. 6, with the distribution of the control parameter $\alpha$ used to calculate this impedance is given in appendix C. As expected the distribution is symmetric about the value of $x=0$ (the transverse position of the line current source), and oscillates as required to convert the incident field into radiation directed along the surface normal.
Figure 6: Impedance tensor distribution for cylindrical to plane wave conversion: The constant amplitude of the reflected plane wave is a factor of 0.09 times the maximum of the incident field at the impedance boundary.
Fig. 7 shows a numerical simulation of the out of plane and in plane components of the electric field for a line current source placed a distance $\lambda/4$ above the impedance distribution shown in Fig. 6. In Fig. 7 (a) we plot the $E_{y}$ electric field component, showing the expected field generated by the line current source. Fig. 7 (b) shows the $E_{x}$ field component, from which we can identify the desired plane wave propagating normal to the impedance boundary. Note that in Fig. 7 (b) we can observe large magnitude features in the fields at the surface that are not part of the desired output field. These maxima are features that depend on the mesh used in the solver and occur where the determinant of the impedance matrix changes sign.
Figure 7: Numerical verification of cylindrical to plane wave conversion: (a) The $\bm{\hat{y}}$ polarised electric field excited by the line current. (b) The $\bm{\hat{x}}$ polarized plane wave resulting from the reflection of the incident field from the impedance boundary.
This example shows the applicability of our method to find passive, tensorial impedance distributions that convert between any specified incident waveform and any desired reflected one.
VI Realization of the metasurface design
The above formalism shows how we can obtain a passive, reciprocal tensorial impedance distribution to realise any given wave transformation. However, in most cases we cannot directly implement an impedance distribution. We must find a unit cell of the metasurface that gives an adequate approximation to the desired impedance. Here we give an example of the procedure that can be used to find an implementation. We take a discrete set of points in the $\lambda\leq x\leq 2\lambda$ section of the impedance distribution shown in Fig. 6. Taking an implementation as an elliptical patch on dielectric layer, tuned near the half wave resonance (see appendix D), we then diagonalise the impedance distribution shown in Fig. 6 by applying a rotation matrix (i.e. choosing the angle of the ellipse as a function of position on the surface). The diagonalised distribution is shown as the blue “target” lines in Fig. 8, and the dimensions of the elliptical patch are varied and related to the surface impedance at a given set of points on the surface as described in e.g.Patel and Grbic (2013); Smith et al. (2024), the results are here shown as the dots in Fig. 8. To obtain the full impedance, each patch is then rotated using the same matrix used to diagonalize the designed impedance matrix.
Note that here we have chosen to ignore the real part of the impedance, which is small in this part of the distribution. To implement the anisotropic loss required by the full design we can e.g. include a set of lossy conducting wires, or anisotropic dielectric in the unit cell. Connected to this, in many cases the real and imaginary parts of the tensorial surface impedance may not be simultaneously diagonalized using a rotation of the unit cell. In these cases we must use a unit cell containing e.g. dielectrics or stacked conducting patches with different principal axes, rotating their relative angle and the angle of the full unit cell to obtain the desired tensorial impedance distribution.
Figure 8: Example realization of impedance profile: Using the elliptical patch unit cell shown in Appendix D, we calculated the eigenvalues, $Z_{xx}^{\prime}$ and $Z_{yy}^{\prime}$ (panels a and b), of the impedance tensor (dots), varying the dimensions of the patch to fit the desired distribution of eigenvalues calculated from Eq. (LABEL:eq:reciprocal_Z) shown in Fig. (6 (solid line). Note the jumps in eigenvalues in panels (a) and (b) simply correspond to the ordering given by the numerical linear algebra routine and do not affect the form of the impedance tensor (the corresponding jump in angle in panel c corrects this change of order). To reproduce the tensor we then determined the rotation angle $\theta_{d}$ necessary to diagonalize the impedance tensor (panel c), thus determining the rotation of each patch on the surface.
VII Summary and Conclusions
@@
Impenetrable metasurfaces are often characterized in terms of a scalar valued surface impedance, the value of which encodes both the phase and amplitude of the reflected field. However, as has been previously recognised, the use of a tensorial surface impedance—practically implemented via an array of anisotropic, polarizable elements—offers a much richer set of possible field transformations. Here we have exploited the non-uniqueness of the metasurface parameters to solve the problem of designing tensorial metasurfaces that enact an arbitrary field transformation while remaining reciprocal, passive, or both.
@@
Our first finding is that it is always possible to find a reciprocal metasurface, whatever the wave transformation we wish to perform. The tensorial surface impedance is given by Eq. (LABEL:eq:reciprocal_Z) and there are an infinite number of such surfaces, parameterized in terms of a single complex number, $\alpha$. In general this family of metasurfaces will exhibit gain and loss for some incident fields, showing that through carefully implementing loss and/or gain within the surface, we may obtain multiplexed functionality. As an example, we showed in Sec. IV that a metasurface rotating $x$ polarized fields by an angle $\theta$ could simultaneously show a variety of different functionalities for yy–polarized radiation, e.g. infinite amplification, absorption, or rotation.
@@
Nevertheless, despite a large body of fascinating work (e.g. Fleury et al. (2014)) and recent experimental progress Tapar and Kishen (2021); Y. et al. (2022), metasurfaces with gain are much more difficult to implement than lossy ones. Our second finding is that, provided the complex parameter $\alpha$ satisfies the inequality (12), we can simultaneously ensure the metasurface is reciprocal and passive (without gain) for arbitrary incident and scattered fields. Surprisingly the surface parameters remain non–unique, and there remain an infinite family of surfaces that perform the same transformation. One important point is that the right hand side of the inequality (12) diverges when the initial design is made $100\%$ efficient (zero dissipation, all energy reflected). This actually reduces the design freedom, implying that the free parameter $\alpha$ must be taken to infinity in order that the surface be both lossless and reciprocal. Therefore, when the efficiency of the design is reduced (e.g. reducing the overall amplitude of the reflected field), we find that this increases the range of reciprocal, passive metasurfaces that can enact the transformation. This was demonstrated in two examples of polarization rotating metasurface, where e.g. Fig. 5b shows that reducing the efficiency immediately yields an increased range of passive reciprocal surfaces that would otherwise require ${\rm Re}[\alpha]\to\infty$. We have thus found that a careful choice of loss within the metasurface actually aids the design, allowing for polarization multiplexed functionality, at the price of reducing the efficiency.
@@ Appendix A Reflection from a tensor impedance surface
Appendix B Obtaining Fields of Line current
In section V we calculated the impedance tensor necessary to reflect a TE polarized cylindrical wave into a TM polarized plane wave. Even though an analytic solution may be available, in general we can find both incident and reflected fields numerically. In our example we modelled the line current in free space using COMSOL multiphysics COMSOL Multiphysics , as shown in Fig. 9.
Figure 9: Determining the incident field: Electric field due to a line current in free space, the central circular region containaing a line current source shifted from the origin in the $z$ direction by $\lambda/4$. The red line indicates where the fields are extracted for use as incident fields in the boundary condition given in Eq. 2 to calculate the distribution shown in Fig. 6
Tabulating the fields along the line $z=\lambda/4$ below the line current source gives us the incident electric field and magnetic fields in Eq.2. These are shown in Fig. 10.
Figure 10: Numerically calculated fields at surface: Electric and magnetic fields evaluated at a distance $\lambda/4$ below a $\bm{\hat{y}}$ polarised line current, indicated as a red line in Fig. 9
Appendix C Obtaining Alpha for spatially varying fields
To calculate the values of the free parameter $\alpha$ appearing in Eq. 12 we write it in the form, $\alpha=C+Re^{i\phi}$ with the radius $R$,
$R=\frac{\sqrt{|a|^{2}-4\Re(adc^{*})+4|c|^{2}\Re(b)}}{2|c|}$
and the centre given by,
$C=-\frac{a^{*}-2dc^{*}}{2|c|}.$
where the variables $a$, $b$, $c$, and $d$ given respectively by;
$a$ $=\bm{\hat{h}}.\bm{\hat{h}}$
$b$ $=e_{1}(\bm{\hat{n}}\times\bm{\hat{h}}^{\star})\cdot\bm{\hat{h}}$
$c$ $=\sqrt{\frac{1}{4{\rm Re}(\bar{e_{2}})}}(\bm{\hat{h}}\cdot(\bm{\hat{n}}\times\bm{\hat{h}}^{*})$
$d$ $=\sqrt{\frac{1}{4{\rm Re}(\bar{e_{2}})}}(\bar{e_{1}^{*}}+e_{1}(\bm{\hat{n}}\times\bm{\hat{h}}^{*})\cdot(\bm{\hat{n}}\times\bm{\hat{h}}^{*}))$ (27)
Through the parameterization in terms of an angle $\phi$ (not related to any physical angle!), a range of impedance distributions can be found, all of which are passive. For the impedance solution given in Fig. 6 a constant value of $\phi=\pi$ is chosen.
Appendix D Example realisation
To approximate the impedance distribution shown in Fig. 8 we used an elliptical patch on a dielectric layer on a grounded copper sheet. The parameters of the unit cell are shown in Fig. 11.
Figure 11: Unit cell design: Unit cell used to approximate the impedance distribution shown in Fig. 8. The top oval patch layer and bottom ground plane layer are copper. The middle dielectric layer is Ventec-VT 6710 ($\epsilon_{r}=10.2+0.024i$)
By changing the major/minor axes of the elliptical patch in $x$ and $y$ (the lattice vectors of the metasurface), we can extract the impedance from either the surface fieldsPatel and Grbic (2013) or numerically computed reflection coefficients Smith et al. (2024). The surface impedance calculated from the reflection coefficients is shown in Fig. 12. The impedance distribution over the full range of patch variations was obtained using FEM COMSOL Multiphysics using 80 steps in both the $0.6\leq D_{x}\leq 0.9$ and $0.6\leq D_{y}\leq 0.9$ ($D_{x}$ and $D_{y}$ being the lengths of the elliptical patch in the $\hat{x}$ and $\hat{y}$ directions respectively). The range was then interpolated to obtain detailed functions of the surface impedance over the range to allow for fitting to the requried surface impedance distribution.
Figure 12: Simulated impedance values: Impedance eigenvalues obtained from finite element modelling of the periodic array of unit-cells shown in Fig. 11 of appendix D, for varying $x$ and $y$ diameter of the ellipse (periodicity of square lattice, $a$). These values are derived for a plane wave at normal incident.